The p?b values for the dibasic base B are p?b1=2.10 and p?b2=7.67. Calculate the pH at each of the points in the titration of 50.0 mL of a 0.80 M B(aq) solution with 0.80 M HCl(aq).

What is the pH before addition of any HCl? pH=

What is the pH after addition of 25.0 mL HCl? pH=

What is the pH after addition of 50.0 mL HCl? pH=

What is the pH after addition of 75.0 mL HCl? pH=

What is the pH after addition of 100.0 mL HCl? pH=
Answer #1

*b = 7.948103 B(OH) ² BloH) + oH ههه kb =213%258 in auxion 0.8-2 - 0.8-150.8 as kb, is less !. 6,352x10-3 - 0163520102 a= 0.7TpH = pky, & hogy 7 [B(OH) F = 2:10 + log on to - 2-10- 14-2.10 = 11.90 some Hel moles = 0.04 moles. equivalence point slottclassmate Date Page B(OH) + Hw - 0.04 B0H). +49 0.04 0.08 0.04 B/OH) ++ H+ oivu oon 0.04 Bt2 + How = a.266 ka, ako Bt² of Hao